An empty rubber balloon has a mass of 12 grams. The balloon is filled with helium at STP. The helium filled balloon has a radius of 50 cm. What is the force of gravity on the balloon?
1
Expert's answer
2019-05-07T09:51:23-0400

The  force of gravity on the balloon is


"W=(m+m_{He})g=(m+\rho_{He} V)g"

"V=\frac{4}{3}\pi r^3"

"W=(m+m_{He})g=(m+\frac{4}{3}\pi r^3\rho_{He} )g"

"W=(0.012+\frac{4}{3}\pi (0.5)^3(0.1785) )9.8=1.0\ N"